OpenStudy (anonymous):

How do I set up this problem? Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart?

OpenStudy (radar):

d=st d (distance) equal s (speed) times t (time) A careful reading of this problem will provide d and s. 45 = 10t + 5t 45= 15t t = 3 hours. Verify: (10)(3) +5(3)=30 + 15 = 45 45 = 45. Note. since they are traveling in opposite directions the distance of 45 is equal to the sum of their respective distance traveled.

OpenStudy (radar):

did you follow with understanding?

OpenStudy (anonymous):

yes

OpenStudy (radar):

Great, and good luck with your studies.